​10. C&A's potato chip filling process has a lower specification limit of 9.5 oz. and an upper specification limit of 10.5 oz.

10. C&A's potato chip filling process has a lower specification limit of 9.5 oz. and an upper specification limit of 10.5 oz. The standard deviation is 0.3 oz. and the mean is 10 oz. What is the process capability index for the chip filling process? 
Answer #1

Process Capability Index(CpK) = Min(CpU, CpL)

CpU = (Upper Specification Limit - Mean)/3*Standard Deviation

= (10.5 - 10)/3*0.3

= 0.5/0.9

= 0.556

CpL = (Mean - Lower Specification Limit)/3*Standard Deviation

= (10 - 9.5)/3*0.3

= 0.556

CpK = Min(0.556, 0.556)

= 0.556